# MAGMA: A Multi-Graph based Agentic Memory Architecture for AI Agents

### v1

@@ Abstract
Memory-Augmented Generation (MAG) extends Large Language Models with external memory to support long-context reasoning, but existing approaches largely rely on semantic similarity over monolithic memory stores, entangling temporal, causal, and entity information. This design limits interpretability and alignment between query intent and retrieved evidence, leading to suboptimal reasoning accuracy. In this paper, we propose MAGMA, a multi-graph agentic memory architecture that represents each memory item across orthogonal semantic, temporal, causal, and entity graphs. MAGMA formulates retrieval as policy-guided traversal over these relational views, enabling query-adaptive selection and structured context construction. By decoupling memory representation from retrieval logic, MAGMA provides transparent reasoning paths and fine-grained control over retrieval. Experiments on LoCoMo and LongMemEval demonstrate that MAGMA consistently outperforms state-of-the-art agentic memory systems in long-horizon reasoning tasks.
@@
Dongming Jiangα, Yi Liα, Guanpeng Liβ and Bingzhe Liα αUniversity of Texas at Dallas βUniversity of Florida {dongming.jiang, yi.li3, bingzhe.li}@utdallas.edu; liguanpeng@ufl.edu
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks Brown2020; achiam2023gpt; Wei2022, yet they remain limited in their ability to maintain and reason over long-term context. These models process information within a finite attention window, and their internal representations do not persist across interactions, causing earlier details to be forgotten once they fall outside the active context Brown2020; Beltagy2020. Even within a single long sequence, attention effectiveness degrades with distance due to attention dilution, positional encoding limitations, and token interference, leading to the well-known “lost-in-the-middle” and context-decay phenomena liu2024lost; Press2021. Moreover, LLMs lack native mechanisms for stable and structured memory, resulting in inconsistent recall, degraded long-horizon reasoning, and limited support for tasks requiring persistent and organized memory Khandelwal2018; maharana2024evaluating.
@@
Despite their promise, current MAG systems exhibit structural and operational limitations that constrain their effectiveness in long-term reasoning li2025memos; chhikara2025mem0; xu2025mem; packer2023memgpt; rasmussen2025zep; wang2025mirix; Kang2025. Most existing approaches store past interactions in monolithic repositories or minimally structured memory buffers, relying primarily on semantic similarity, recency, or heuristic scoring to retrieve relevant content. For example, A-Mem xu2025mem organizes past interactions into Zettelkasten-like memory units that are incrementally linked and refined, yet their retrieval pipelines rely primarily on semantic embedding similarity, missing the relations such as temporal or causal relationships. Cognitive-inspired frameworks like Nemori nan2025nemori introduce principled episodic segmentation and representation alignment, enabling agents to detect event boundaries and construct higher-level semantic summaries. However, their memory structures are still narrative and undifferentiated, with no explicit modeling of distinct relational dimensions.
@@
    We demonstrate that MAGMA consistently outperforms state-of-the-art agentic memory systems on long-context benchmarks including LoCoMo and LongMemEval, while reducing retrieval latency and token consumption relative to prior systems. The code is open-sourced [^1] https://github.com/FredJiang0324/MAMGA.
@@ 2 Background
Existing Large Language Models (LLMs) face fundamental challenges in handling long-term agentic interactions. These challenges stem from the inherent limitations of fixed-length contexts, which result in fragmented memory and an inability to maintain narrative coherence over time. The evolution of long-term consistency in LLMs is shifted from Context-Window Extension (Beltagy2020; Press2021; kang2025lm2; qian2025memorag), Retrieval-Augmented Generation (RAG) (lewis2020retrieval; jiang2025rago; wang2024m; jiang2024longrag; gutierrez2025rag; lin2025cache) to Memory-Augmented Generation (MAG).
Retrieval-oriented approaches enrich the model with an external, dynamic memory library, giving rise to the paradigm of Memory-Augmented Generation (MAG) zhong2024memorybank; park2023generative; huang2024emotional. Formally, unlike static RAG, MAG maintains a time-variant memory $\mathcal{M}_{t}$ that evolves via a feedback loop:
@@
Some prior schemes focused on structuring the intermediate states or relationships of memory to enable better reasoning. Think-in-Memory (TiM) (liu2023think) stores evolving chains-of-thought to maintain consistency. A-MEM (xu2025mem) draws inspiration from the Zettelkasten method, organizing knowledge into an interconnected note network. More recently, graph-based approaches like GraphRAG (edge2024graphrag) and Zep (rasmussen2025zep) structure memory into knowledge graphs to capture cross-document dependencies. We provide a detailed discussion of related work in Appendix A.
However, prior work typically organizes memory around associative proximity (e.g., semantic similarity) rather than mechanistic dependency (kiciman2023causal). As a result, such methods can retrieve what occurred but struggle to reason about why, since they lack explicit representations of causal structure, leading to reduced accuracy in complex reasoning tasks (jin2023cladder; zhang2025igniting).
@@ 3.2 Data Structure Layer
where $c_{i}$ denotes the event content (e.g., observations, actions, or state changes), $\tau_{i}$ is a discrete timestamp anchoring the event in time, and $\mathbf{v}_{i}\in\mathbb{R}^{d}$ is a dense representation indexed in the vector database (johnson2019billion). The attribute set $\mathcal{A}_{i}$ captures structured metadata such as entity references, temporal cues, or contextual descriptors, enabling hybrid retrieval that integrates semantic similarity with symbolic and structural constraints.
@@ 3.3 Query Process: Adaptive Hierarchical Retrieval
Stage 2 - Multi-Signal Anchor Identification: Before initiating graph traversal, the system first identifies a set of anchor nodes that serve as entry points into the memory graph. To ensure robustness across query modalities, we fuse signals from dense semantic retrieval, lexical keyword matching, and temporal filtering using Reciprocal Rank Fusion (RRF) Cormack2009:
@@ 3.4 Memory Evolution (Write and Update)
Long-term reasoning requires not only effective retrieval, but also a memory substrate that can adapt and reorganize as experience accumulates. MAGMA addresses this requirement through a structured memory evolution scheme that incrementally refines its multi-relational graph over time. Specifically, the transition from $\mathcal{G}_{t}$ to $\mathcal{G}_{t+1}$ is governed by a dual-stream process that decouples latency-sensitive ingestion from compute-intensive consolidation Kumaran2016, balancing short-term responsiveness with long-term reasoning fidelity.
@@ 4.1 Experimental Setup
Datasets. We evaluate long-term conversational capability using two widely adopted benchmarks: (1) LoCoMo (maharana2024evaluating): which contains ultra-long conversations (average length of 9K tokens) designed to assess long-range temporal and causal retrieval. (2) LongMemEval (wu2024longmemeval): a large-scale stress-test benchmark with an average context length exceeding 100K tokens, used to evaluate scalability and memory retention stability over extended interaction horizons..
@@
  • A-MEM (xu2025mem): A biological-inspired, self-evolving memory system that dynamically organizes agent experiences.
  • Nemori (nan2025nemori): A graph-based memory utilizing a "predict-calibrate" mechanism for episodic segmentation.
  • MemoryOS(Kang2025) : A semantic-focused memory operating system employing a hierarchical storage strategy.
Metrics. Following standard evaluation protocols, we primarily use the LLM-as-a-Judge score (zheng2023judging) to assess the accuracy of different methods. The detailed evaluation prompt used for the judge model is provided in the appendix. For completeness, we also report token-level F1 and BLEU-1 (papineni-etal-2002-bleu).
@@ 4.5 Ablation Study
First, removing Traversal Policy results in the largest performance drop, with the Judge score decreasing from 0.700 to 0.637. This confirms that intent-aware routing is critical: without it, retrieval degenerates into a static graph walk that introduces structurally irrelevant information and degrades reasoning quality. Second, removing either Causal Links or the Temporal Backbone leads to comparable and substantial performance losses (0.644 and 0.647, respectively), indicating that causal structure and temporal ordering provide complementary, non-substitutable axes of reasoning. Finally, removing Entity Links causes a smaller but consistent decline (0.700 to 0.666), highlighting their role in maintaining entity permanence and reducing hallucinations in entity-centric queries.
@@ 6 Limitations
While MAGMA demonstrates strong empirical performance, it has several limitations. First, the quality of the constructed memory graph depends on the reasoning fidelity of the underlying Large Language Models used during asynchronous consolidation. This dependency is a shared limitation of agentic memory systems that rely on LLM-based structural inference, as they are susceptible to extraction errors and hallucinations (pan2024unifying; xi2025rise; wadhwa2023revisiting). Although MAGMA employs structured prompts and conservative inference thresholds to reduce spurious links, erroneous or missing relations may still arise and propagate to downstream retrieval. Nevertheless, our experimental results indicate that, even under these constraints, agentic memory systems such as MAGMA substantially outperform traditional baselines, including full-context approaches, in long-horizon reasoning tasks.
@@
Finally, most existing agentic memory systems, including MAGMA, are primarily evaluated on long-context conversational and agentic benchmarks such as LoCoMo and LongMemEval. While these benchmarks effectively stress temporal and causal reasoning, they do not cover the full range of settings in which agentic memory may be required(hu2025memory). Extending MAGMA to other scenarios, such as multimodal agents or environments with heterogeneous observation streams, may require additional adaptation and calibration. Addressing these broader evaluation settings remains an important research direction for future work.
@@ Appendix A Related Work
Context-window Extension. A direct line of work extends the effective context length of Transformers by modifying attention or positional extrapolation. Longformer beltagy2020longformer introduces sparse attention patterns to scale to long documents, reducing quadratic cost while retaining locality and selected global connectivity. ALiBi press2021train (Attention with Linear Biases) enables length extrapolation by injecting distance-aware linear biases into attention scores, improving robustness when testing on longer sequences than those seen in training. Recent efforts also add explicit memory modules or hybrid mechanisms to push beyond pure attention-window scaling. For example, LM2 kang2025lm2 proposes a decoder-only architecture augmented with an auxiliary memory to mitigate long-context limitations. MemoRAG qian2025memorag similarly emphasizes global-memory-enhanced retrieval to boost long-context processing when raw context is insufficient or inefficient. While these approaches improve long-range coverage, they do not, by themselves, address the continual, evolving, and write-back nature of agent memory required for multi-session interactions.
@@
Retrieval Augmented Generation. RAG lewis2020retrieval augments an LLM with external retrieval over a fixed corpus, classically retrieving supporting passages and conditioning generation on them. Subsequent work explores better integration with long-context models and more scalable retrieval pipelines. LongRAG jiang2024longrag studies how to exploit long-context LLMs together with retrieval, improving the ability to incorporate larger retrieved evidence sets. Other systems focus on structuring the retrieved memory space or optimizing the RAG serving stack: M-RAG wang2024m uses multiple partitions to encourage fine-grained retrieval focus, while RAGO jiang2025rago provides a systematic framework for performance optimization in RAG serving. However, standard RAG typically assumes a static knowledge base. In contrast, agentic settings require memory that is continuously updated (the feedback loop described in the main text). This motivates the shift to MAG systems, where memory is dynamic and evolves with interaction histories.
@@
Memory Augmented Generation and Agent Memory Systems. MAG systems maintain and update an external memory over time, enabling agents to accumulate knowledge, preserve identity, and remain coherent across sessions. Early and representative directions include memory construction and write-back strategies for long-term agent behavior, such as MemoryBank zhong2024memorybank and generative agents style architectures that emphasize persistent profiles and evolving state grounded in past interactions nan2025nemori; maharana2024evaluating. A growing body of work adopts systems metaphors and designs: MemGPT packer2023memgpt frames LLM agents with an operating-system-like memory hierarchy, emphasizing paging and controlled context management. More recent memory OS systems propose explicit storage hierarchies and controllers (e.g., MemoryOS kang2025memory, MemOS li2025memos) to manage persistence and retrieval policies at scale. In addition, practical agent-memory stacks (e.g., Zep rasmussen2025zep) offer temporal knowledge-graph-based memory services aimed at real-world deployment constraints.
@@
Structured memory: chains-of-thought and graph-based representations. Beyond flat text buffers or vector stores, several methods explicitly structure memory to support reasoning. Think-in-Memory (TiM) stores evolving chains-of-thought to improve consistency across long-horizon reasoning, while A-MEM xu2025mem is inspired by Zettelkasten-style linking of notes/experiences. These methods highlight the value of representing intermediate reasoning traces or explicit links, but many retrieval pipelines still predominantly rely on semantic similarity as the primary access mechanism. Graph-based approaches have recently gained traction as a way to capture cross-document and cross-episode dependencies. GraphRAG edge2024local builds entity-centric graphs and community summaries to answer more global questions over large corpora. Zep proposes a temporally-aware knowledge-graph engine (Graphiti) that synthesizes conversational and structured business data while preserving historical relations. The main text notes these graph-based lines explicitly and motivates a key gap: many systems organize memory around associative proximity (semantic relatedness) rather than mechanistic dependency.
@@
Causal reasoning and long-horizon evaluation. Causal reasoning has been highlighted as both important and challenging for LLMs. The work kiciman2023causal study LLMs’ ability to generate causal arguments across multiple causal tasks and emphasize robustness/failure modes, reinforcing that what happened retrieval is not sufficient for why reasoning in many settings. Benchmarking efforts such as LoCoMo maharana2024evaluating stress long-range temporal and causal dynamics in multi-session conversations and provide evaluation tasks that expose long-horizon memory deficits. The paper’s experimental setup also uses LongMemEval wu2024longmemeval as an ultra-long context stress test, and evaluates via LLM-as-a-Judge protocols standard in modern instruction-following evaluation. Overall, prior work demonstrates steady progress in (i) scaling context length, (ii) improving retrieval pipelines, and (iii) building structured, evolving memories for agents. The main text positions MAGMA within this trajectory by explicitly targeting multi-relational structure (semantic/temporal/causal/entity) and intent-aware retrieval control.

### v2

@@ Abstract
Memory-Augmented Generation (MAG) extends Large Language Models with external memory to support long-context reasoning, but existing approaches largely rely on semantic similarity over monolithic memory stores, entangling temporal, causal, and entity information. This design limits interpretability and alignment between query intent and retrieved evidence, leading to suboptimal reasoning accuracy. In this paper, we propose MAGMA, a multi-graph agentic memory architecture that represents each memory item across orthogonal semantic, temporal, causal, and entity graphs. MAGMA formulates retrieval as policy-guided traversal over these relational views, enabling query-adaptive selection and structured context construction. By decoupling memory representation from retrieval logic, MAGMA provides transparent reasoning paths and fine-grained control over retrieval. Experiments on LoCoMo and LongMemEval demonstrate that MAGMA consistently outperforms state-of-the-art agentic memory systems in long-horizon reasoning tasks. The open-source code is available at: https://github.com/FredJiang0324/MAGMA.
@@
Dongming Jiangα, Yi Liα, Guanpeng Liβ, Bingzhe Liαthanks: Corresponding author αDepartment of Computer Science, The University of Texas at Dallas βDepartment of Electrical and Computer Engineering, University of Florida {dongming.jiang, yi.li3, bingzhe.li}@utdallas.edu liguanpeng@ufl.edu
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks Brown et al. (2020); Achiam et al. (2023); Wei et al. (2022), yet they remain limited in their ability to maintain and reason over long-term context. These models process information within a finite attention window, and their internal representations do not persist across interactions, causing earlier details to be forgotten once they fall outside the active context Brown et al. (2020); Beltagy et al. (2020a). Even within a single long sequence, attention effectiveness degrades with distance due to attention dilution, positional encoding limitations, and token interference, leading to the well-known “lost-in-the-middle” and context-decay phenomena Liu et al. (2024); Press et al. (2021a). Moreover, LLMs lack native mechanisms for stable and structured memory, resulting in inconsistent recall, degraded long-horizon reasoning, and limited support for tasks requiring persistent and organized memory Khandelwal et al. (2018); Maharana et al. (2024).
@@
Despite their promise, current MAG systems exhibit structural and operational limitations that constrain their effectiveness in long-term reasoning Li et al. (2025); Chhikara et al. (2025); Xu et al. (2025); Packer et al. (2023); Rasmussen et al. (2025); Wang and Chen (2025); Kang et al. (2025a). Most existing approaches store past interactions in monolithic repositories or minimally structured memory buffers, relying primarily on semantic similarity, recency, or heuristic scoring to retrieve relevant content. For example, A-Mem Xu et al. (2025) organizes past interactions into Zettelkasten-like memory units that are incrementally linked and refined, yet their retrieval pipelines rely primarily on semantic embedding similarity, missing the relations such as temporal or causal relationships. Cognitive-inspired frameworks like Nemori Nan et al. (2025) introduce principled episodic segmentation and representation alignment, enabling agents to detect event boundaries and construct higher-level semantic summaries. However, their memory structures are still narrative and undifferentiated, with no explicit modeling of distinct relational dimensions.
@@
    We demonstrate that MAGMA consistently outperforms state-of-the-art agentic memory systems on long-context benchmarks including LoCoMo and LongMemEval, while reducing retrieval latency and token consumption relative to prior systems. The code is open-sourced [^1] https://github.com/FredJiang0324/MAGMA.
@@ 2 Background
Existing Large Language Models (LLMs) face fundamental challenges in handling long-term agentic interactions. These challenges stem from the inherent limitations of fixed-length contexts, which result in fragmented memory and an inability to maintain narrative coherence over time. The evolution of long-term consistency in LLMs is shifted from Context-Window Extension (Beltagy et al., 2020a; Press et al., 2021a; Kang et al., 2025c; Qian et al., 2025), Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Jiang et al., 2025; Wang et al., 2024; Jiang et al., 2024; Gutiérrez et al., 2025; Lin et al., 2025) to Memory-Augmented Generation (MAG).
Retrieval-oriented approaches enrich the model with an external, dynamic memory library, giving rise to the paradigm of Memory-Augmented Generation (MAG) Zhong et al. (2024); Park et al. (2023); Huang et al. (2024). Formally, unlike static RAG, MAG maintains a time-variant memory $\mathcal{M}_{t}$ that evolves via a feedback loop:
@@
Some prior schemes focused on structuring the intermediate states or relationships of memory to enable better reasoning. Think-in-Memory (TiM) (Liu et al., 2023) stores evolving chains-of-thought to maintain consistency. A-MEM (Xu et al., 2025) draws inspiration from the Zettelkasten method, organizing knowledge into an interconnected note network. More recently, graph-based approaches like GraphRAG (Edge et al., 2024a) and Zep (Rasmussen et al., 2025) structure memory into knowledge graphs to capture cross-document dependencies. We provide a detailed discussion of related work in Appendix A.
However, prior work typically organizes memory around associative proximity (e.g., semantic similarity) rather than mechanistic dependency (Kiciman et al., 2023). As a result, such methods can retrieve what occurred but struggle to reason about why, since they lack explicit representations of causal structure, leading to reduced accuracy in complex reasoning tasks (Jin et al., 2023; Zhang et al., 2025).
@@ 3.2 Data Structure Layer
where $c_{i}$ denotes the event content (e.g., observations, actions, or state changes), $\tau_{i}$ is a discrete timestamp anchoring the event in time, and $\mathbf{v}_{i}\in\mathbb{R}^{d}$ is a dense representation indexed in the vector database (Johnson et al., 2019). The attribute set $\mathcal{A}_{i}$ captures structured metadata such as entity references, temporal cues, or contextual descriptors, enabling hybrid retrieval that integrates semantic similarity with symbolic and structural constraints.
@@ 3.3 Query Process: Adaptive Hierarchical Retrieval
Stage 2 - Multi-Signal Anchor Identification: Before initiating graph traversal, the system first identifies a set of anchor nodes that serve as entry points into the memory graph. To ensure robustness across query modalities, we fuse signals from dense semantic retrieval, lexical keyword matching, and temporal filtering using Reciprocal Rank Fusion (RRF) Cormack et al. (2009):
@@ 3.4 Memory Evolution (Write and Update)
Long-term reasoning requires not only effective retrieval, but also a memory substrate that can adapt and reorganize as experience accumulates. MAGMA addresses this requirement through a structured memory evolution scheme that incrementally refines its multi-relational graph over time. Specifically, the transition from $\mathcal{G}_{t}$ to $\mathcal{G}_{t+1}$ is governed by a dual-stream process that decouples latency-sensitive ingestion from compute-intensive consolidation Kumaran et al. (2016), balancing short-term responsiveness with long-term reasoning fidelity.
@@ 4.1 Experimental Setup
Datasets. We evaluate long-term conversational capability using two widely adopted benchmarks: (1) LoCoMo (Maharana et al., 2024): which contains ultra-long conversations (average length of 9K tokens) designed to assess long-range temporal and causal retrieval. (2) LongMemEval (Wu et al., 2024): a large-scale stress-test benchmark with an average context length exceeding 100K tokens, used to evaluate scalability and memory retention stability over extended interaction horizons..
@@
  • A-MEM (Xu et al., 2025): A biological-inspired, self-evolving memory system that dynamically organizes agent experiences.
  • Nemori (Nan et al., 2025): A graph-based memory utilizing a "predict-calibrate" mechanism for episodic segmentation.
  • MemoryOS(Kang et al., 2025a) : A semantic-focused memory operating system employing a hierarchical storage strategy.
Metrics. Following standard evaluation protocols, we primarily use the LLM-as-a-Judge score (Zheng et al., 2023) to assess the accuracy of different methods. The detailed evaluation prompt used for the judge model is provided in the appendix. For completeness, we also report token-level F1 and BLEU-1 (Papineni et al., 2002).
@@ 4.4 System Efficiency Analysis
Table 5: Single-graph ablation study on LoCoMo.
Graph Configuration Multi-Hop Temporal Open-Domain Single-Hop Adversarial Overall
Causal Only 0.470 0.460 0.430 0.650 0.680 0.590
Temporal Only 0.440 0.620 0.450 0.650 0.520 0.577
Entity Only 0.485 0.420 0.460 0.640 0.450 0.531
Full MAGMA 0.528 0.650 0.517 0.776 0.742 0.700
@@ 4.5 Ablation Study
First, removing the Adaptive Policy results in the largest performance drop, with the Judge score decreasing from 0.700 to 0.637. This confirms that intent-aware routing is critical: without it, retrieval degenerates into a static graph walk that introduces structurally irrelevant information and degrades reasoning quality. Second, removing either Causal Links or the Temporal Backbone leads to comparable and substantial performance losses (0.644 and 0.647, respectively), indicating that causal structure and temporal ordering provide complementary, non-substitutable axes of reasoning. Finally, removing Entity Links causes a smaller but consistent decline (0.700 to 0.666), highlighting their role in maintaining entity permanence and reducing hallucinations in entity-centric queries.
To further isolate the contribution of each relation type, we additionally conducted a single-graph-only ablation on LoCoMo, shown in Table 5. The results are consistent with the leave-one-out findings in Table 4. Among single-graph variants, Causal Only achieves the highest overall score (0.590), suggesting that causal relations provide strong logical filtering and robustness against distractor noise. Temporal Only performs best on temporal questions (0.620), confirming that explicit temporal structure is particularly important for sequential reasoning. In contrast, Entity Only obtains the lowest overall score (0.531): while it remains helpful for concept bridging in multi-hop reasoning, it lacks both timeline awareness and logical filtering, which explains why removing entity links causes the smallest drop in the full-system ablation.
Overall, these results show that no single relation type is sufficient to recover MAGMA’s full reasoning capability, as all single-graph variants remain below 0.60 overall. By explicitly decoupling causal, temporal, and entity relations and combining them with adaptive traversal, MAGMA leverages their complementary strengths to achieve the best overall performance.
@@ 6 Limitations
While MAGMA demonstrates strong empirical performance, it has several limitations. First, the quality of the constructed memory graph depends on the reasoning fidelity of the underlying Large Language Models used during asynchronous consolidation. This dependency is a shared limitation of agentic memory systems that rely on LLM-based structural inference, as they are susceptible to extraction errors and hallucinations (Pan et al., 2024; Xi et al., 2025; Wadhwa et al., 2023). Although MAGMA employs structured prompts and conservative inference thresholds to reduce spurious links, erroneous or missing relations may still arise and propagate to downstream retrieval. Nevertheless, our experimental results indicate that, even under these constraints, agentic memory systems such as MAGMA substantially outperform traditional baselines, including full-context approaches, in long-horizon reasoning tasks.
@@
Finally, most existing agentic memory systems, including MAGMA, are primarily evaluated on long-context conversational and agentic benchmarks such as LoCoMo and LongMemEval. While these benchmarks effectively stress temporal and causal reasoning, they do not cover the full range of settings in which agentic memory may be required(Hu et al., 2025; Jiang et al., 2026). Extending MAGMA to other scenarios, such as multimodal agents or environments with heterogeneous observation streams, may require additional adaptation and calibration. Addressing these broader evaluation settings remains an important research direction for future work.
7 Acknowledgment
This work was partially supported by NSF 2343863, 2413520, 2417747 and 2440611. Any opinions, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the NSF.
@@ Appendix A Related Work
Context-window Extension. A direct line of work extends the effective context length of Transformers by modifying attention or positional extrapolation. Longformer Beltagy et al. (2020b) introduces sparse attention patterns to scale to long documents, reducing quadratic cost while retaining locality and selected global connectivity. ALiBi Press et al. (2021b) (Attention with Linear Biases) enables length extrapolation by injecting distance-aware linear biases into attention scores, improving robustness when testing on longer sequences than those seen in training. Recent efforts also add explicit memory modules or hybrid mechanisms to push beyond pure attention-window scaling. For example, LM2 Kang et al. (2025c) proposes a decoder-only architecture augmented with an auxiliary memory to mitigate long-context limitations. MemoRAG Qian et al. (2025) similarly emphasizes global-memory-enhanced retrieval to boost long-context processing when raw context is insufficient or inefficient. While these approaches improve long-range coverage, they do not, by themselves, address the continual, evolving, and write-back nature of agent memory required for multi-session interactions.
@@
Retrieval Augmented Generation. RAG Lewis et al. (2020) augments an LLM with external retrieval over a fixed corpus, classically retrieving supporting passages and conditioning generation on them. Subsequent work explores better integration with long-context models and more scalable retrieval pipelines. LongRAG Jiang et al. (2024) studies how to exploit long-context LLMs together with retrieval, improving the ability to incorporate larger retrieved evidence sets. Other systems focus on structuring the retrieved memory space or optimizing the RAG serving stack: M-RAG Wang et al. (2024) uses multiple partitions to encourage fine-grained retrieval focus, while RAGO Jiang et al. (2025) provides a systematic framework for performance optimization in RAG serving. Furthermore, to improve the reliability of retrieved context, approaches like Self-Correcting RAG Xu et al. (2026) enhance generation faithfulness via selective context filtering and inference-guided search. However, standard RAG typically assumes a static knowledge base. In contrast, agentic settings require memory that is continuously updated (the feedback loop described in the main text). This motivates the shift to MAG systems, where memory is dynamic and evolves with interaction histories.
@@
Memory Augmented Generation and Agent Memory Systems. MAG systems maintain and update an external memory over time, enabling agents to accumulate knowledge, preserve identity, and remain coherent across sessions. Early and representative directions include memory construction and write-back strategies for long-term agent behavior, such as MemoryBank Zhong et al. (2024) and generative agents style architectures that emphasize persistent profiles and evolving state grounded in past interactions Nan et al. (2025); Maharana et al. (2024); Zeng et al. (2025). A growing body of work adopts systems metaphors and designs: MemGPT Packer et al. (2023) frames LLM agents with an operating-system-like memory hierarchy, emphasizing paging and controlled context management. More recent memory OS systems propose explicit storage hierarchies, efficient memory modules, and controllers (e.g., MemoryOS Kang et al. (2025b), MemOS Li et al. (2025), and Hippocampus Li et al. (2026)) to manage persistence, scalability, and retrieval policies at scale. In addition, practical agent-memory stacks (e.g., Zep Rasmussen et al. (2025)) offer temporal knowledge-graph-based memory services aimed at real-world deployment constraints.
@@
Structured memory: chains-of-thought and graph-based representations. Beyond flat text buffers or vector stores, several methods explicitly structure memory to support reasoning. Think-in-Memory (TiM) stores evolving chains-of-thought to improve consistency across long-horizon reasoning, while A-MEM Xu et al. (2025) is inspired by Zettelkasten-style linking of notes/experiences. These methods highlight the value of representing intermediate reasoning traces or explicit links, but many retrieval pipelines still predominantly rely on semantic similarity as the primary access mechanism. Graph-based approaches have recently gained traction as a way to capture cross-document and cross-episode dependencies. GraphRAG Edge et al. (2024b) builds entity-centric graphs and community summaries to answer more global questions over large corpora. Zep proposes a temporally-aware knowledge-graph engine (Graphiti) that synthesizes conversational and structured business data while preserving historical relations. The main text notes these graph-based lines explicitly and motivates a key gap: many systems organize memory around associative proximity (semantic relatedness) rather than mechanistic dependency.
@@
Causal reasoning and long-horizon evaluation. Causal reasoning has been highlighted as both important and challenging for LLMs. The work Kiciman et al. (2023) study LLMs’ ability to generate causal arguments across multiple causal tasks and emphasize robustness/failure modes, reinforcing that what happened retrieval is not sufficient for why reasoning in many settings. Benchmarking efforts such as LoCoMo Maharana et al. (2024) stress long-range temporal and causal dynamics in multi-session conversations and provide evaluation tasks that expose long-horizon memory deficits. The paper’s experimental setup also uses LongMemEval Wu et al. (2024) as an ultra-long context stress test, and evaluates via LLM-as-a-Judge protocols standard in modern instruction-following evaluation. Overall, prior work demonstrates steady progress in (i) scaling context length, (ii) improving retrieval pipelines, and (iii) building structured, evolving memories for agents. The main text positions MAGMA within this trajectory by explicitly targeting multi-relational structure (semantic/temporal/causal/entity) and intent-aware retrieval control.